Of the 130 students who took a math exam, 90 correctly answered the first question , 62 correctly answered the second question, and 50 correctly answered both questions.
a. How many students correctly answered either the first or second question?
b. How many students did not answer either of the two questions correctly?
c. How many students answered either the fist or second question correctly, but not both?

Answer by stanbon(48558)

Of the 130 students who took a math exam, 90 correctly answered the first question , 62 correctly answered the second question, and 50 correctly answered both questions.
a. How many students correctly answered either the first or second question?
#= 90 + 62 - 50 = 102
--------
b. How many students did not answer either of the two questions correctly?
# = 130-102 = 28
----------

c. How many students answered either the fist or second question correctly, but not both?
# = 90-50 + 62 - 50 = 52
--------
Cheers,
Stan H.